OpenStudy (anonymous):

Prove or Disprove: Given k,n ∈N If k|n! then k|i for some 0 ≤ i ≤ n.

ganeshie8 (ganeshie8):

At last a question that I can answer !

ganeshie8 (ganeshie8):

Counterexample : let k = 6, n = 3

OpenStudy (anonymous):

6|3!=6 but what is i here?

ganeshie8 (ganeshie8):

since n = 3, i is 0, 1, 2, 3

ganeshie8 (ganeshie8):

Hey wait, every integer divides 0, so are you sure the question has no typoes ?

OpenStudy (anonymous):

yes it's right

ganeshie8 (ganeshie8):

Then the given statement is trivially true just let \(i=0\)

ganeshie8 (ganeshie8):

Every integer divides 0 so every k | 0

OpenStudy (anonymous):

maybe proof by induction then

ganeshie8 (ganeshie8):

I still feel the question has a typo

OpenStudy (anonymous):

no i'm sure is right

ganeshie8 (ganeshie8):

\(0\le i \le n\) makes no sense perhaps it should be \(0\color{red}{\lt} i \le n\) ?